A farmer who has fixed amounts of land and capital finds that total product is 24 for the first worker hired, 32 when two workers are hired, 37 when three are hired, and 40 when four are hired. The farmer's product sells for $3 per unit, and the wage rate is $13 per worker. How many workers should the farmer hire?

Multiple Choice

  • 1

  • 2

  • 3

  • 4
Answer #1

Answer : The answer is option C.

For 1st worker total cost (TC) = $13

Total revenue (TR) from 1st worker = Price * Quantity = 3 * 24 = $72

Profit from 1st worker = TR - TC = 72 - 13 = $59

For 2 workers, TC = 2 * 13 = $26

TR from 2 workers = 3 * 32 = $96

Profit from 2 workers = TR - TC = 96 - 26 = $70

For 3 workers, TC = 3 * 13 = $39

TR from 3 workers = Price * Quantity = 3 * 37 = $111

Profit from 3 workers = TR - TC = 111 - 39 = $72.

For 4 workers, TC = 4 * 13 = $52

TR from 4 workers = 3 * 40 = $120

Profit from 4 workers = 120 - 52 = $68

As by hiring 3 workers the farmer gets maximum profit of $72, hence the farmer should hire 3 workers. Therefore, option C is correct.